# Robust Market Interventions

### v1

@@ 2.3. A locally linear oligopoly: Market states and signals
The matrix $\bm{D}=\bm{D}(\bm{p}^{0})$ and the status quo quantities $\bm{q}^{0}=\bm{q}(\bm{p}^{0})$ together determine the first-order effects of an intervention $\bm{\sigma}$ on consumer surplus $C$, producer surplus $P$, and total surplus $W$, i.e., the derivatives $\dot{C}_{\bm{\sigma}}$, $\dot{P}_{\bm{\sigma}}$, and $\dot{W}_{\bm{\sigma}}$.
We note that $\bm{D}$ satisfies the following property (Nocke and Schutz, 2017).
Property NSD.
The following assumption is without loss of generality—it holds by suitably adjusting the units of the numeraire $m$, and will be useful later.
Assumption 3.
The quantity vector’s Euclidean norm $\|\bm{q}^{0}\|$ is at most 1.
Fact 1 implies that if the authority knows the pair $(\bm{D},\bm{q}^{0})$ characterizing the market conditions at the status quo, then it knows the local surplus outcomes of interventions. We, therefore, call the tuple $(\bm{D},\bm{q}^{0})$, where $\bm{D}$ is a negative semidefinite matrix with diagonal entries $-1$ and $\bm{q}^{0}$ is a vector of norm at most $1$, the market state.
2.3.1. The authority’s signal
The authority receives a signal about the market state $\bm{\theta}=(\bm{D},\bm{q})$. This signal consists of random variables
[…]
@@ 2.3.1. The authority’s signal
We make the following assumptions about the authority’s signal. Let $\|\bm{E}\|$ denote the spectral norm of a matrix $\bm{E}$, which for a symmetric matrix is equal to its largest eigenvalue. Fix a sequence $b(n)$ and a positive constant $\overline{V}$.
@@
The first part of the assumption bounds the matrix norm of the errors in estimating the normalized Slutsky matrix. In Appendix C, we provide a simple procedure for sampling market data independently across product pairs $(i,j)$ under which this assumption holds with $b(n)=\gamma n^{1/2}$ (for a positive constant $\gamma>0$). We can think of this example as generating $E_{ij}$ that are essentially independent, with variance of constant order (i.e., neither growing or decaying with $n$). This shows that the assumption can hold even when there is no entry of $\bm{D}$ that can be recovered accurately.
@@
If $b(n)=\gamma n^{\beta}$ with $\beta\in(1/2,1)$, there are error structures consistent with 4 where some terms in $\widehat{\bm{D}}$ have large covariances. For example, spatially correlated errors, with sufficiently “distant” parts of the matrix being at most slightly correlated, would satisfy this assumption. [^12] The idea is the same as that of ergodicity-type conditions in time-series settings. An example of a situation that would cause part (1) of 4 to be violated is the entries of $\bm{E}$ all having correlation bounded away from zero (e.g., coming from a common shock to measured complementarities).
@@
The second part of the assumption requires independence of errors across different products’ quantities. The purpose of the assumption is to apply a law of large numbers for estimating outcomes such as the average quantity, as well as various linear combinations of quantities that are important for intervention outcomes. Independence is stronger than we need for our main result stated in 1, and is made to facilitate exposition. When we use this assumption in the proof our main result, 1, we rely on a substantially weaker condition (which is more technical to state) that 4(2) implies (see Section A.2). Concerning the assumption on the norm of $\bm{\varepsilon}$, recall that we assume that $\|\bm{q}^{0}\|\leq 1$; the assumption on $\bm{\varepsilon}$ scales the error to be of the same order of magnitude as the quantity vector.
We maintain Assumptions 1–4 unless specifically noted otherwise.
2.4. Robust interventions
In the linear oligopoly environment, 1 establishes that to determine the first-order surplus effects of any intervention $\bm{\sigma}$, the only additional data needed is the tuple $(\bm{D},\bm{q}^{0})$, which we call the market state and denote by $\bm{\theta}$. We take a prior-free approach to modeling the authority’s uncertainty over $\bm{\theta}$. That is, we assume the authority knows only that $\bm{\theta}$ belongs to some set $\bm{\Theta}$ and does not have a probability distribution over $\bm{\Theta}$. However, the authority knows the distribution of the signal $\bm{t}=(\widehat{\bm{D}},\widehat{\bm{q}}^{0})$ given $\bm{\theta}\in\bm{\Theta}$. We denote this distribution by $\mu_{\bm{\theta}}$. We denote by $\mathcal{T}$ the set of possible values that the signal $\bm{t}$ can take.
Given the authority’s signal $\bm{t}$, the authority designs an intervention rule [^13] This should be measurable in a suitable sense, which is clear in our application.
$\bm{R}:\mathcal{T}\to\bm{\Sigma},$
prescribing an intervention $\bm{\sigma}\in\bm{\Sigma}$ for every possible signal $\bm{t}$.
A market outcome is a tuple $(\bm{\theta},\bm{\sigma})$ consisting of a market state and an intervention. We call this pair the outcome because it determines production, consumption and transfers. A property is a subset of the set $\bm{\Theta}\times\bm{\Sigma}$ of all possible outcomes. The authority wants to design an intervention rule that robustly achieves a desired property. We are interested in understanding which properties can be achieved with high probability in all market states that the authority considers possible:
Definition 1.
An intervention rule $\bm{R}$ achieves a property $\bm{X}$ $\epsilon$–robustly if the following holds: For every $\bm{\theta}\in\bm{\Theta}$, we have
$\mu_{\bm{\theta}}\left(\{\bm{t}:(\bm{\theta},\bm{R}(\bm{t}))\in\bm{X})\}\right)\geq 1-\epsilon.$
The only randomness in the definition is in the draw of the signal.
Notice that the definitions we have made do not depend on any details of the linear oligopoly environment. In any environment where $(\bm{\theta},\bm{\sigma})$ fully determines all outcomes of interest to the authority and there is a known distribution over signals given states, the definition can be applied.
3. Recoverable structure
We define a new concept, which we call recoverable structure, that plays a key role in the analysis of robust interventions that we will present in Section 4. Recoverable structure imposes conditions on the pattern of complements and substitutes (which we will call interactions) among products, as summarized by the Slutsky matrix. It is a condition ensuring that some latent pattern of product interactions can be inferred robustly through the noisy observation and, at the same time, this latent structure is correlated with market quantities. We now introduce it formally and provide an economic interpretation.
A vector in $\mathbb{R}^{n}$ describes a bundle of products. Given $\bm{D}$, we are interested in the subspace of bundles spanned by eigenvectors of $\bm{D}$ with large eigenvalues. Formally, let $\mathcal{L}(\bm{D},M)\subseteq\mathbb{R}^{n}$ be the subspace of the bundle space that is spanned by the eigenvectors of $\bm{D}$ with eigenvalues at least $M$ in absolute value.
Definition 2.
To understand the definition, note that a Slutsky matrix, by virtue of being symmetric, can be orthogonally diagonalized: it can be written as a linear combination of orthogonal rank-one matrices:
Here, $\lambda_{1},\lambda_{2},\ldots,\lambda_{n}$ are the eigenvalues of $\bm{D}$ (which are nonpositive numbers because $\bm{D}$ is negative semidefinite), ordered from greatest to least in absolute value, and $(\bm{u}^{1},\dots,\bm{u}^{n})$ is a corresponding basis of orthonormal eigenvectors. All the summands are orthogonal to each other, and $|\lambda_{\ell}|$ is the norm of the contribution of the corresponding summand.
Having market states with $(M,\delta)$-recoverable structure means that (i) $\bm{D}$ has eigenvectors with eigenvalues larger than $M$ in absolute value, so the first summation in the above formula is nonzero, and furthermore (ii) the vectors $\bm{u}^{\ell}$ in that sum can jointly account for a significant portion of the status quo quantities. As we will formally explain later, if we set $M$ to be larger than the (suitably measured) size of the noise $\bm{E}$ in the observation of $\bm{D}$, then condition (i) will ensure that the low-dimensional subspace $\mathcal{L}(\bm{D},M)$ can be estimated precisely despite $\bm{D}$ being observed with noise; it is intuitive that this is valuable for intervening in the market. However, if quantities were nearly orthogonal to this subspace, then it turns out that the recoverable structure would only make a small difference to surplus outcomes; the requirement (ii) entailed in the definition rules this out by imposing that the quantities in the market have nontrivial projection onto the recoverable subspace. In this sense the definition states that $(\bm{D},\bm{q}^{0})$ has a structure which can be recovered and is relevant to the market.
We provide an example to illustrate the notion of recoverable structure.

### v2

@@ 3.2.2. Application to the stochastic block model
In the stochastic block model, if $\gamma(n)n^{1/2}\to\infty$, then there is a $\delta>0$ such that, for large enough $n$, the set of market states has $(M(n),\delta)$-recoverable structure for some constant $\delta>0$ and a sequence $M(n)$ with $M(n)/n^{1/2}\to\infty$.
The intuition is straightforward: because there are only finitely many types, some blocks in $\bm{D}$ must be large; as in Example 1, a large block with entries of order $\gamma(n)$ gives rise to an eigenvalue of magnitude $\gamma(n)n$. The proof also shows that the projection onto the corresponding eigenspace of $\bm{q}^{0}$ arising from a generic $\bar{\bm{q}}^{0}$ has norm bounded below by a number independent of $n$.
To see how the presence of aggregate structure relates to large-scale complementarities in the example, recall that we assumed $\bar{\bm{D}}$ itself satisfies NSD, and so its diagonal entries are $-1$. This means that there is a large block on the diagonal of $\bm{D}$ with negative entries of magnitude $\gamma(n)$—large-scale complementarities just as in Example 1.
Thus, we can now keep the stochastic block model in mind as a canonical example of recoverable structure with $M(n)\gg\sqrt{n}$.
3.3. A key tool: Spectral price theory
When recoverable structure is present, information about the top eigenvectors of $\bm{D}$ will be recoverable with high precision—under suitable assumptions on errors—using standard statistical results. We now describe how an authority can use information solely about top eigenvectors to design an intervention that increases surplus robustly.
The key tool is a decomposition of surplus pass-through in spectral terms, building on a spectral description of the pass-through of an intervention to prices and quantities. Denoting by $\bm{U}$ the matrix whose $\ell^{\text{th}}$ column is the $\ell^{\text{th}}$ eigenvector $\bm{u}^{\ell}$ of $\bm{D}$, and by $\bm{\Lambda}$ the matrix whose non-diagonal elements are zero and whose $\ell^{\text{th}}$ diagonal element is $\lambda_{\ell}$, we have:
$\bm{D}=\bm{U}\bm{\Lambda}\bm{U}^{\mathsf{T}}.$
An intervention $\bm{\sigma}$ that subsidizes (or taxes) a single product will in general affect not only the prices and quantities of that product but also those of other products, whose equilibrium values are all connected through strategic interactions. If we think of the eigenvectors $\bm{u}^{\ell}$ as representing bundles, then these bundles have the important property that an intervention $\bm{\sigma}\propto\bm{u}^{\ell}$ in the direction of such a bundle will only affect the price $\bm{u}^{\ell}\cdot\bm{p}$ and quantity $\bm{u}^{\ell}\cdot\bm{q}$ of that bundle, leaving the prices and quantities of the bundles corresponding to the other eigenvectors unchanged. Generally, we can decompose $\bm{\sigma}=\sum_{\ell}(\bm{u}^{\ell}\cdot\bm{\sigma})\bm{u}^{\ell}$ into a combination of $n$ orthogonal interventions, each in the direction of an eigenvector. We can use this decomposition to solve the oligopoly game and obtain simple expressions for the pass-through of the intervention in terms of the eigenvalues of $\bm{D}$.
Lemma 2.
Proof.
Thus, we can study the price and quantity pass-throughs of each of these $n$ interventions separately across eigenvectors. In particular, each unit of subsidy in direction $\bm{u}^{\ell}$ exclusively passes through to the price and quantities of bundle $\bm{u}^{\ell}$, and it does so with coefficients $-(1+|\lambda_{\ell}|)^{-1}$ and $|\lambda_{\ell}|(1+|\lambda_{\ell}|)^{-1}$, respectively.
Note that the magnitudes of the price and quantity pass-throughs in the different $\bm{u}^{\ell}$ are ordered according to their corresponding eigenvalues: The larger is $|\lambda_{\ell}|$, the less a given subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$ reduces prices, but the more it increases quantities. This asymmetry is the result of two opposing forces: On the one hand, the strategic interactions among firms imply that the equilibrium price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ is less sensitive to the subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$ the larger is $|\lambda_{\ell}|$. On the other, the demand $\bm{u}^{\ell}\cdot\bm{q}^{0}$ is more sensitive to the price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ the larger [^18] Indeed, it follows from (4) that $\bm{U}^{\mathsf{T}}\bm{q}^{0}=\bm{\Lambda}\bm{U}^{\mathsf{T}}\bm{p}^{0}$, so the slope of the demand $\bm{u}^{\ell}\cdot\bm{q}^{0}$ with respect to own price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ is equal to $\lambda_{\ell}$. is $|\lambda_{\ell}|$; this is just a fact about the market’s demand function, rather than equilibrium pricing. 2 shows that the second effect dominates the first in the sense that the larger is $|\lambda_{\ell}|$, the more sensitive is the equilibrium quantity $\bm{u}^{\ell}\cdot\bm{q}^{0}$ to the subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$.
Lemma 3.
The pass-throughs to consumer, producer, and total surpluses are:
Proof.
The effect of the intervention on consumer surplus is $\dot{C}=-\bm{q}^{0}\cdot\dot{\bm{p}}$. Multiplying this equation by $\bm{U}\bm{U}^{\mathsf{T}}$ gives $\dot{C}=-\bm{U}^{\mathsf{T}}\bm{q}^{0}\cdot\bm{U}^{\mathsf{T}}\dot{\bm{p}}$. Similarly, the effect of the intervention on producer surplus is $\dot{P}=\bm{q}^{0}\cdot(\dot{\bm{p}}+\bm{\sigma})+(\bm{p}-\bm{c})\cdot\dot{\bm{q}}=2\bm{q}^{0}\cdot\dot{\bm{q}}$. Multiplying the equation by $\bm{U}\bm{U}^{\mathsf{T}}$ yields the expression of $\dot{P}$ in the Lemma. The expression for $\dot{W}$ is obtained by aggregating $\dot{P}$, $\dot{C}$ and the intervention expenditure. ∎
3 shows that the effect of an intervention on consumer, producer, or total surplus is a weighted sum of pass-throughs to each of the eigenvectors $\bm{u}^{\ell}$—with the weight being the corresponding bundle’s quantity.
We now use 3 to illustrate how—by setting $\bm{\sigma}$ equal to $\bm{u}^{1}$—the authority may achieve the highest total surplus per dollar spent possible subject to the constraint that the change in consumer surplus is not negative (recall 2). Because $\bm{u}^{1}$ is orthogonal to all the other $\bm{u}^{\ell}$ with $\ell\neq 1$, such intervention changes only the price and quantity of the bundle $\bm{u}^{1}$:
leading to an overall change in consumer and producer surplus equal to
Recoverable structure requires that $|\lambda_{1}|\to\infty$ with $n$, and so (by inspection of the equations) the interventions will achieve $\dot{C}/\dot{S}\to 0$ and $\dot{P}/\dot{S}\to 2$.
4. Conditions for robust surplus improvements
In 1, we reported a complete information benchmark that shows what is possible with no observation errors. In 2, we provided an example in which a natural constrained-efficient outcome among these is achievable under considerable observation errors. This section presents our main result, which generalizes this example.
4.1. Assumptions on errors
@@
We now add to our maintained assumptions the following assumption about the authority’s signal. Let $\|\bm{E}\|$ denote the spectral norm of a matrix $\bm{E}$, which for a symmetric matrix is equal to its largest eigenvalue. Fix a sequence $b(n)$ and a positive constant $\overline{V}$.
@@
The first part of the assumption bounds the matrix norm of the errors in estimating the normalized Slutsky matrix. In Appendix C, we provide a simple procedure for sampling market data independently across product pairs $(i,j)$ under which this assumption holds with $b(n)=\gamma n^{1/2}$ (for a positive constant $\gamma>0$). We can think of this example as generating $E_{ij}$ that are essentially independent, with variance of constant order (i.e., neither growing nor decaying with $n$). This shows that the assumption can hold even when there is no entry of $\bm{D}$ that can be recovered accurately.
@@
If $b(n)=\gamma n^{\beta}$ with $\beta\in(1/2,1)$, there are error structures consistent with 4 where some terms in $\widehat{\bm{D}}$ have large covariances. For example, spatially correlated errors, with sufficiently “distant” parts of the matrix being at most slightly correlated, would satisfy this assumption. [^19] The idea is analogous to ergodicity-type conditions in time-series settings. An example that would violate part (1) of 4 is the entries of $\bm{E}$ all having correlation bounded away from zero (e.g., arising from a common shock to measured complementarities).
@@
The second part of the assumption requires independence of errors across different products’ quantities. The purpose of the assumption is to apply a law of large numbers for estimating outcomes such as the average quantity, as well as various linear combinations of quantities that are important for intervention outcomes. Independence is stronger than we need for our main result stated in 1, and is made to facilitate exposition. When we use this assumption in the proof of our main result, 1, we rely on a substantially weaker but more technical condition that 4(2) implies (see Section A.3). Regarding the assumption on the norm of $\bm{\varepsilon}$, recall that we assume that $\|\bm{q}^{0}\|\leq 1$; the assumption on $\bm{\varepsilon}$ scales the error to be of the same order of magnitude as the quantity vector.
4.2. Main result
Our main result, 1, examines what can be implemented robustly when the authority has partial information and the economy has a recoverable structure that is strong enough relative to the noise. It shows that interventions exist that robustly protect consumers from surplus loss and implement market surplus equal to the upper bound achievable by an omniscient authority (the upper bound given by 1(2)).
Theorem 1.
Let $M(n)$ be an increasing sequence with $M(n)/b(n)\to_{n}\infty$, and fix $\delta>0$. Assume the set of market states has $(M(n),\delta)$-recoverable structure. For every $\epsilon>0$ and for every target expenditure $s>0$, the following properties can be simultaneously achieved $\epsilon$–robustly for sufficiently large $n$:
  1. (i)
    The sum of marginal consumer and producer surplus gains $\dot{C}+\dot{P}$ is at least twice the marginal expenditure, up to a small multiplicative error: $\dot{C}+\dot{P}\geq(2-\epsilon)\dot{S}$.
  2. (ii)
  3. (iii)
    The marginal expenditure is arbitrarily close to the target expenditure $s$, i.e., $|\dot{S}-s|<\epsilon$.
The condition in 1 stipulates that the market has $(M(n),\delta)$-recoverable structure for some $M(n)$ that asymptotically dominates $b(n)$. This lower bound on $M(n)$ ensures that the recoverable structure in $\bm{D}$ is substantially larger than the norm of the error matrix $\bm{E}$, which is $O(b(n))$ under 4. Note also that the assumptions on errors are satisfied by the stochastic block model of Section 3.1, so 2 is a direct corollary of 1.202020This follows since the entries of $\bm{E}$ were taken to be independent (Dallaporta, 2012).
Point (ii) states that it is possible to achieve this while leaving all households’ welfare essentially unchanged; indeed, under the policy we construct, producers fully capture the surplus gains. Note that by 1, the welfare gain described in point (i) is essentially the maximum total surplus change that an omniscient authority could implement with the same expenditure without reducing consumer surplus.
Our notion of $\epsilon$–robustness means that these surplus properties are achieved ex post with high probability. A natural question is whether they are also achieved in expectation (since, in principle, realizations with very negative surplus could occur with low probability). In our setting, it turns out that all the analysis would be unaffected if we added good ex ante expected performance to the definition of robustness. [^22] This is because the surplus pass-throughs are supported on $[0,1]$ and all quantities in the proofs are bounded. Thus, convergence in probability is equivalent to convergence in $L^{1}$, and so our proofs extend to show close approximations to the omniscient benchmark in terms of ex ante surplus.
To prove 1, we apply a statistical method that accurately identifies a subspace spanned by top eigenvectors of the Slutsky matrix from noisy observations. We then show that interventions projecting exclusively onto these recoverable subspaces possess the desirable welfare properties stated in 1.
4.3. Recovering and using the subspace of top eigenvectors
[…]
4.3.1. The Davis–Kahan theorem
Recall that if $M(n)\gg b(n)$, then $(M(n),\delta)$-recoverable structure requires that the normalized Slutsky matrix $\bm{D}$ has eigenvalues that, in absolute value, are much larger than $b(n)$; we will refer to such eigenvalues simply as “large” from now on.
The key tool in our statistical exercise that leverages this assumption is the Davis–Kahan theorem. Under the hypothesis that some eigenvalues of $\bm{D}$ are large, this theorem guarantees that, despite the noise in $\bm{E}$, the large eigenvalues of the observed matrix are good approximations of the true large eigenvalues. In other words, the noise in $\bm{E}$ cannot cause the large eigenvalues of $\bm{D}$ to become “mixed up” with the eigenvalues far away in the spectrum; see Figure 1 for an illustration. The theorem also permits the recovery of eigenvectors. More precisely, this theorem has the following two central implications in our setting:
  1. (i)
  2. (ii)
    the eigenvectors of $\widehat{\bm{D}}$ associated with such eigenvalues are in the span of eigenvectors of $\bm{D}$ with large eigenvalues, i.e., the eigenvectors of $\widehat{\bm{D}}$ associated with such eigenvalues can be expressed (up to a small error) as linear combinations of eigenvectors of $\bm{D}$ with large eigenvalues.
This powerful result forms the core of our strategy to recover and use structure underlying the oligopoly demand robustly under noise. To facilitate the illustration, we impose a stronger assumption on $\bm{D}$: that the largest eigenvalue of $\bm{D}$ is sufficiently well-separated from all other eigenvalues by a “gap” much larger than $b(n)$. Under this condition, the Davis–Kahan Theorem yields an even stronger implication: we can use $\widehat{\bm{D}}$ to recover a normalized eigenvector $\widehat{\bm{u}}^{1}$ that correlates almost perfectly with the corresponding eigenvector [^23] In Section 5 we develop Example 2 to illustrate our main result. Figure 2 illustrates the similarity between the true $\bm{u}^{1}$ (in panel B) and the estimated $\widehat{\bm{u}}^{1}$ (panel D) for a case where this stronger “gap” condition holds. $\bm{u}^{1}$.
To use this, observe that 2 and 3 together imply
while expenditure is
Let us use the recovered $\widehat{\bm{u}}^{1}$ to design an intervention with $\bm{\sigma}\propto\widehat{\bm{u}}^{1}$. The Davis–Kahan Theorem allows us to treat $\widehat{\bm{u}}^{1}$ as effectively equal to its true counterpart $\bm{u}^{1}$ with a very small error, so from now on we will treat $\bm{u}^{1}$ as known. By choosing the sign of $\bm{\sigma}$ to ensure $(\bm{u}^{1}\cdot\bm{q}^{0})(\bm{u}^{1}\cdot\bm{\sigma})$ is positive. Then we can see from the equations above that $\dot{W}$ will closely approximate $\dot{S}$, since $\lambda_{1}\gg b(n)$ implies $\frac{|\lambda_{1}|}{1+|\lambda_{1}|}\approx 1$. Moreover, if we know $\bm{u}^{1}\cdot\bm{q}^{0}$ and this differs from zero (which is a requirement of recoverable structure), we can scale the intervention to be of the size that we desire, and achieve $\dot{S}=s$.
This argument contains some wishful thinking, however. When we arranged the sign of $\bm{\sigma}$ so that $(\bm{u}^{1}\cdot\bm{q}^{0})(\bm{u}^{1}\cdot\bm{\sigma})$ is positive, we did not consider that we have only a noisy observation $\widehat{\bm{q}}^{0}$ of $\bm{q}^{0}$. So part of the challenge of the proof is to manage the observation error that makes $\widehat{\bm{q}}^{0}$ different from ${\bm{q}}^{0}$, and to show that we can obtain a correct estimate of the sign with probability tending to $1$ as $n\to\infty$. If we fail to do this correctly, our intervention actually decreases efficiency with positive probability. This explains the need for the second part of Assumption 4 on the error $\bm{\varepsilon}$ in the quantity signal.
However, bounded noise alone is insufficient: if $\bm{u}^{1}\cdot\bm{q}^{0}$ is very small, there may be no hope for consistently recovering the true magnitude or sign of $\bm{u}^{1}\cdot\bm{q}^{0}$ from the signal $\widehat{\bm{u}}^{1}\cdot(\bm{q}^{0}+\bm{\varepsilon})$ even with well-behaved noise: the asymptotically small noise could still overwhelm a similarly decaying underlying mean $\bm{u}^{1}\cdot\bm{q}^{0}$. Such an unrecoverability would make it impossible to orient and scale our intervention appropriately. The definition of recoverable structure prevents this problem by requiring that the projection of $\bm{q}^{0}$ onto eigenvectors with large eigenvalues is bounded away from zero.
The special case of our main result that this discussion makes plausible is: If the largest eigenvalue of $\bm{D}$ is well-separated from others and if $\bm{q}^{0}\cdot\bm{u}^{1}$ is not vanishingly small, then a subsidy profile proportional to $\bm{u}^{1}$ can, if it is suitably scaled, achieve all the properties of 1.
4.3.3. The more general result
The proof of the main result improves on this sketch in two ways. First, it does not rely only on the eigenspace spanned by $\bm{u}^{1}$. Instead, it uses a potentially much larger eigenspace of $\widehat{\bm{D}}$. The general intervention projects $\widehat{\bm{q}}^{0}$ onto $\mathcal{L}(\bm{D},M(n))$, the eigenspace of all eigenvectors of $\bm{D}$ with eigenvalues larger than $M(n)$. This makes it easier for the analog of $\bm{u}^{1}\cdot\bm{q}^{0}$ not to be too small, since the projection of $\bm{q}^{0}$ onto a larger eigenspace will have a larger norm. Second, the general proof dispenses with assuming that any eigenvalues are well-separated. Instead, it handles any possible spectrum of $\bm{D}$ subject to our maintained assumptions. This introduces considerable complexity, as it is no longer generally possible to recover any true eigenvector $\bm{u}^{\ell}$ with any accuracy. We instead work directly with a recovered eigenspace that generalizes the span of $\widehat{\bm{u}}^{1}$. We show that despite limited knowledge of individual true eigenvectors of $\bm{D}$ underlying this space, we can use the fact that all of them have large eigenvalues to generalize our argument for showing that (10) and (11) can be made very close and nonzero with a feasible intervention. This is where the arguments go beyond standard applications of the Davis–Kahan theorem.
5. Illustration: Interventions based on noisy demand estimates
We begin this section by providing an example illustrating the notion of recoverable structure, and then use this example to demonstrate the effects of the intervention rule that taxes and subsidizes firms in the direction of the eigenvector associated with the largest eigenvalue of the observed Slutsky matrix.